# Leave No Context Behind: Efficient Infinite Context Transformers with Infini-attention

### 10 Apr 2024 v1

@@ 2.1 Infini-attention
As shown Figure 1, our Infini-attention computes both local and global context states and combine them for its output. Similar to multi-head attention (MHA), it maintains $H$ number of parallel compressive memory per attention layer ($H$ is the number of attention heads) in addition to the dot-product attention.
@@ 2.1.1 Scaled Dot-product Attention
For MHA, we compute $H$ number of attention context vectors for each sequence element in parallel, concatenate them along the second dimension and then finally project the concatenated vector to the model space to obtain attention the output.
@@ 2.2 Memory and Effective Context Window
Our Infini-Transformer enables an unbounded context window with a bounded memory footprint. To illustrate this, Table 1 lists the previous segment-level memory models with their context-memory footprint and effective context length defined in terms of model parameters and input segment length. Infini-Transformer has a constant memory complexity of $d_{key}\times d_{value}+d_{key}$ for storing compressed context in $M_{s}$ and $z_{s}$ for each head in single layer while for the other models, the complexity grows along with the sequence dimension - the memory complexity depends either on the cache size for Transformer-XL (Dai et al., 2019), Compressive Transformer (Rae et al., 2019) and Memorizing Transformers (Wu et al., 2022) or on the soft-prompt size for RTM (Bulatov et al., 2022) and AutoCompressors (Ge et al., 2023).
@@ 4 Related Work
Compressed input representations can be viewed as a summary of past sequence segments (Rae et al., 2019; Chevalier et al., 2023). Along this direction, more recent works have been utilizing a Transformer LLM itself to compress input sequence for efficient long-context modeling (Bulatov et al., 2022; Chevalier et al., 2023; Ge et al., 2023; Mu et al., 2024). However, the previous segment-level compression methods, including Compressive Transformers (Rae et al., 2019) still discard the memory entries of old segments in order to free up space for the new ones, limiting their context window to the most recent segments. This is in contrast to our Infini-attention that computes incremental memory updates to a fixed amount of memory parameters in a recurrent fashion.
Long-context continual pre-training. There is a line of work that extends the do-product attention layers and continues to train LLMs for long-context (Xiong et al., 2023; Fu et al., 2024). The attention extensions include incorporating sparsity into the attention layer (Chen et al., 2023b; Ratner et al., 2022; Mohtashami & Jaggi, 2024) as well as manipulating the position encodings (Chen et al., 2023a; Peng et al., 2023) Although the position encoding-based methods such as position interpolation techniques (Chen et al., 2023a) can be data efficient as they only adjust the positional bias in the attention layer, they are still costly for inference.
@@
Efficient attention. The efficient attention techniques attempt to improve the efficiency of the dot-product attention with an approximation or a system-level optimization. Multiple directions have been explored for different forms of efficient attention approximation, including sparsity-based (Child et al., 2019; Beltagy et al., 2020; Sukhbaatar et al., 2021; Ding et al., 2023) and linear attention approximation (Shen et al., 2018; Katharopoulos et al., 2020; Schlag et al., 2021). Among those, the linear attention variants are closely related to the associative memory matrix (Schlag et al., 2020; 2021) and the metalearned neural memory (Munkhdalai et al., 2019), where KV bindings (Smolensky, 1990) are stored in Fast-Weights (Hinton & Plaut, 1987; Schmidhuber, 1992; Ba et al., 2016) that are modified in with respect to new contextual information. More recently, system-level optimization techniques have been proposed by leveraging specific hardware architecture to make the exact attention computation more efficient (Dao et al., 2022; Liu et al., 2023).

### 09 Aug 2024 v2

@@ 1 Introduction
2 Background
Recurrent Neural Networks (RNNs) process a single token $x_{t}$ at each step $t$ and computes a recurrent hidden state $h_{t}$ to represent an entire input sequence (Hochreiter & Schmidhuber, 1997; Maass et al., 2002):
$h_{t}=RNN(x_{t},h_{t-1}).$ (1)
The RNN computation is very efficient since the model maintains only a fixed-size vector $h_{t}$ for input sequence. However, for processing long sequences it becomes difficult to store entire contextual information into a single fixed-size vector and this limitation had implications on RNNs utility in certain tasks (Kaiser & Sutskever, 2015). To address the limitation, people extended the standard RNNs with an external memory component that can be read from and written to. One such an instance is Metalearned Neural Memory (MNM) (Munkhdalai et al., 2019):
$h_{t},\theta_{t}=MNM(x_{t},h_{t-1},\theta_{t-1}).$ (2)
MNM learns an additional memory state $\theta$ parameterized by a feed-forward neural network (FFN) and uses query, key and value vectors (QKV) to interact with the memory, similar to the attention mechanism. To store information, it modifies the parameters of the FFN by using the key vectors as input and the value vectors for the target, and to read memory entries, it forward-passes the query vectors through the memory FFN and retrieves its corresponding value. Like RNNs, the memory state is still bounded in MNM.
Unlike the RNNs, the attention mechanism however doesn’t maintain a recurrent state and only performs a feed-forward computation on input sequence segment $X_{s}$:
$O_{s}=attention(X_{s}).$ (3)
The attention output $O_{s}$ is simply passed to the next layer and no state is carried over to the next input sequence $X_{s+1}$} at the same attention layer. In the attention layer, in order to capture the dependency between the consequent segments $X_{s}$ and $X_{s+1}$, one needs to process them altogether at the same time and this process becomes a bottleneck requiring large computational resources as the length of input sequence grows more and more. To improve the efficiency while still being able to benefit from the expressiveness of the attention mechanism, this work introduces a recurrent attention layer.
@@ 3.1 Infini-attention
As shown Figure 1, our Infini-attention is a recurrent attention mechanism that computes both local and global context states and combine them for its output. Similar to multi-head attention (MHA), it maintains $H$ number of parallel compressive memory per attention layer ($H$ is the number of attention heads) in addition to the dot-product attention and like the RNNs and MNM, it maintains a recurrent memory state to efficiently track the long sequence context:
$O_{s},M_{s}=infini\text{-}attention(X_{s},M_{s-1})$ (4)
@@ 3.1.1 Scaled Dot-product Attention
For MHA, we compute $H$ number of attention context vectors for each sequence element in parallel, concatenate them along the second dimension and then finally project the concatenated vector to the model space to obtain the attention output.
@@ 3.2 Memory and Effective Context Window
Our Infini-Transformer enables an unbounded context window with a bounded memory footprint. To illustrate this, Table 1 lists the previous segment-level memory models with their context-memory footprint and effective context length defined in terms of model parameters and input segment length. Infini-Transformer has a constant memory complexity of $d_{key}\times d_{value}+d_{key}$ for storing compressed context in $M_{s}$ and $z_{s}$ for each head in single layer while for the other models, the complexity grows along with the sequence dimension - the memory complexity depends either on the cache size for Transformer-XL (Dai et al., 2019), Compressive Transformer (Rae et al., 2019) and Memorizing Transformers (Wu et al., 2022) or on the soft-prompt size for RMT (Bulatov et al., 2022) and AutoCompressors (Ge et al., 2023).
@@ 4 Experiments
4.1 Implementation details
Segment chunking. we forward-pass the entire input text a Transformer model and then perform segment chunking at each Infini-attention layer - in this way, perform a minimal modification to the existing Transformer implementation. The Infini-attention layer segments the input and process it segment by segment and concatenates back the segments to pass the original-length segment as output to the next layer.
Back-propagation through time (BPTT). Each Infini-attention layer is trained with back-propagation through time (Werbos, 1988) by computing the gradient w.r.t the compressive memory states, similar to how RNNs are trained. To save memory, we perform gradient checkpoint when processing the sequence segment by segment.
Position Embeddings (PE). As shown Figure 1, we don’t use position embeddings for the key and query vectors of the compressive memory to store only global contextual information in the long-term memory. The PEs were applied to the QK vectors only after the compressive memory reading and update.
@@ 5 Related Work
Compressed input representations can be viewed as a summary of past sequence segments (Rae et al., 2019; Chevalier et al., 2023). Along this direction, more recent works have been utilizing a Transformer LLM itself to compress input sequence for efficient long-context modeling (Bulatov et al., 2022; Chevalier et al., 2023; Ge et al., 2023; Mu et al., 2024; Hwang et al., 2024). However, the previous segment-level compression methods, including Compressive Transformers (Rae et al., 2019) still discard the memory entries of old segments in order to free up space for the new ones, limiting their context window to the most recent segments. This is in contrast to our Infini-attention that computes incremental memory updates to a fixed amount of memory parameters in a recurrent fashion.
Long-context continual pre-training. There is a line of work that extends the dot-product attention layers and continues to train LLMs for long-context (Xiong et al., 2023; Fu et al., 2024). The attention extensions include incorporating sparsity into the attention layer (Chen et al., 2023b; Ratner et al., 2022; Mohtashami & Jaggi, 2024) as well as manipulating the position encodings (Chen et al., 2023a; Peng et al., 2023). Although the position encoding-based methods such as position interpolation techniques (Chen et al., 2023a) can be data efficient as they only adjust the positional bias in the attention layer, they are still costly for inference.
@@
Efficient attention. The efficient attention techniques attempt to improve the efficiency of the dot-product attention with an approximation or a system-level optimization. Multiple directions have been explored for different forms of efficient attention approximation, including sparsity-based (Child et al., 2019; Beltagy et al., 2020; Sukhbaatar et al., 2021; Ding et al., 2023; Xiao et al., 2024) and linear attention approximation (Shen et al., 2018; Katharopoulos et al., 2020; Schlag et al., 2021). Among those, the linear attention variants are closely related to the associative memory matrix (Schlag et al., 2020; 2021) and the metalearned neural memory (Munkhdalai et al., 2019), where KV bindings (Smolensky, 1990) are stored in Fast-Weights (Hinton & Plaut, 1987; Schmidhuber, 1992; Ba et al., 2016) that are modified in with respect to new contextual information. More recently, system-level optimization techniques have been proposed by leveraging specific hardware architecture to make the exact attention computation more efficient (Dao et al., 2022; Liu et al., 2023).
@@ 6 Conclusion
Acknowledgments
We would like to thank Dongseong Hwang for their help implementing efficient sequence unrolling mechanism with the jax scan function. We would also like to thank Aditya Gupta, Kalpesh Krishna, Tu Vu and Alexandra Chronopoulou for their feedback.